# Information Discovery in e-Commerce

### v1

@@ Appendix A Datasets
In this appendix, we list benchmark datasets that are relevant for studying information discovery in e-commerce. We follow the topical of our chapters, and divide the datasets into $5$ types: e-commerce infrastructures, e-commerce user modeling, e-commerce search, e-commerce recommendation, and e-commerce QA & Dialogues.
@@ A.1 Datasets for e-commerce infrastructures
  • Taobao short title dataset [sun2018multi]: This dataset contains $411,246$ title-product pairs in $94$ categories. Each item in the dataset is represented as $\langle Q,K,S\rangle$, where $Q$ denotes the products’ original titles, $K$ refers to the background knowledge about the products, and $S$ represents the human-written short titles.
@@
  • Walmart e-commerce product dataset [mukherjee2021unsupervised]: The dataset contains 5 parts: D-search includes the top 12 million product search queries on Walmart.com and their frequencies over a 1 year period. D-product includes $250,000$ top-selling Walmart products over a 6 month period. D-com-human includes $40,445$ human-generated title compressions from the Walmart catalog across 8 different product categories. D-meta-auto contains $40,000$ meta-training examples. And D-meta-human is a dataset consisting of $16,000$ human-generated 1-shot title compression examples.
@@ A.2 Datasets for e-commerce user modeling
  • Taobao Tianchi consumer dataset [kim2021deep]: The dataset includes responses of users to advertisements of inventory in the user profile and advertising information. The time length of the data is 8 days, and the dataset is divided into four tables: advertisement features, user profiles, past shopping behavior that users engaged in, and who received the advertisement with responses. [^1] https://tianchi.aliyun.com/dataset/dataDetail?dataId=56
@@
  • Feeds user dataset [yi2021debiasedrec]: The feeds dataset is collected on Microsoft News App from August 1, 2020, to September 1, 2020. It contains $643,177$ news items, over $10,000$ users, $320,925$ impressions, and $970,846$ clicks.
@@
  • Twitter user behavior dataset [al2012homophily]: Each attribute dataset consists of approximately 400 labeled Twitter users, 200 with one label (e.g., “female”) and 200 with a second label (e.g., “male”). In addition, all of the friends of these labeled users are identified; for each of these labeled and neighbor users, the most recent 1000 tweets generated by the user were collected.
@@ A.4 Datasets for e-commerce recommendations
  • AliExpress dataset [ahmed2021deep]: This dataset is collected from an online retailer service owned by the Alibaba group. There are about $2,260,923$ records from AliExpress, the data for about fourteen months from January 1, 2019 to February 23, 2020. The dataset contains $1506,850$ users that submitted the reviews against $49,221$ items in 205 different categories, such as electronic, entertainment, education, house, and garden, etc., and the items are rated from 1 to 5 scale.
@@ A.5 Datasets for e-commerce QA and dialogues
  • JDDC e-commerce dialogue dataset [chen2019jddc]: JDDC is a large-scale real scenario Chinese E-commerce conversation corpus, with more than 1 million multi-turn dialogues, 20 million utterances, and 150 million words, which contains conversations about after-sales topics between users and customer service staffs in an e-commerce scenario. JDDC was updated with multi-modal customer service information in 2021 [zhao2021jddc].
  • E-commerce Dialogue Corpus dataset [zhang2018modeling]: The dataset is collected from the real-world conversations between customers and customer service staff on Taobao. It contains over 5 types of conversations (i.e., commodity consultation, logistics express, recommendation, negotiation, and chitchat) based on over 20 commodities. [^11] https://drive.google.com/file/d/154J-neBo20ABtSmJDvm7DK0eTuieAuvw/view?usp=sharing
Table A.5 summarizes the key statistics for the datasets listed abvoe.
@@
JDDC e-commerce dialogue dataset 1,024,196 20,451,337 150,716,172 7.4 20 [chen2019jddc]

### v3

@@ Appendix A Datasets
In this appendix, we list benchmark datasets that are relevant for studying information discovery in e-commerce. We follow the topical organization of our sections, and divide the datasets into five types: e-commerce infrastructures, e-commerce user modeling, e-commerce search, e-commerce recommendation, and e-commerce QA & dialogues.
@@ A.1 Datasets for e-commerce infrastructures
  • Taobao short title dataset [sun2018multi]: This dataset contains 411,246 title-product pairs in 94 categories. Each item in the dataset is represented as a triple $\langle Q,K,S\rangle$, where $Q$ denotes the products’ original titles, $K$ refers to the background knowledge about the products, and $S$ represents the human-written short titles.
@@
  • Walmart e-commerce product dataset [mukherjee2021unsupervised]: The dataset contains five parts: D-search includes the top 12 million product search queries on Walmart.com and their frequencies over a one year period. D-product includes 250,000 top-selling Walmart products over a six month period. D-com-human includes 40,445 human-generated title compressions from the Walmart catalog across eight different product categories. D-meta-auto contains 40,000 meta-training examples. And D-meta-human is a dataset consisting of 16,000 human-generated 1-shot title compression examples.
@@ A.2 Datasets for e-commerce user modeling
  • Taobao Tianchi consumer dataset [kim2021deep]: The dataset includes responses of users to advertisements of inventory in the user profile and advertising information. The time length of the data is eight days, and the dataset is divided into four tables: advertisement features, user profiles, past shopping behavior that users engaged in, and who received the advertisement with responses. [^1] https://tianchi.aliyun.com/dataset/dataDetail?dataId=56
@@
  • Feeds user dataset [yi2021debiasedrec]: The feeds dataset is collected on Microsoft News App from August 1, 2020 to September 1, 2020. It contains 643,177 news items, over 10,000 users, 320,925 impressions, and 970,846 clicks.
@@
  • Twitter user behavior dataset [al2012homophily]: Each attribute dataset consists of approximately 400 labeled Twitter users, 200 with one label (e.g., “female”) and 200 with a second label (e.g., “male”). In addition, all of the friends of these labeled users are identified; for each of these labeled and neighbor users, the most recent 1,000 tweets generated by the user were collected.
@@ A.4 Datasets for e-commerce recommendations
  • AliExpress dataset [ahmed2021deep]: This dataset is collected from an online retailer service owned by the Alibaba group. There are about 2,260,923 records from AliExpress, the data for about fourteen months from January 1, 2019 to February 23, 2020. The dataset contains 1,506,850 users that submitted reviews against 49,221 items in 205 different categories, such as electronics, entertainment, education, house, and garden, etc., and the items are rated from 1 to 5 scale.
@@ A.5 Datasets for e-commerce QA and dialogues
  • JDDC e-commerce dialogue dataset [chen2020jddc]: JDDC is a large-scale real scenario Chinese E-commerce conversation corpus, with more than one million multi-turn dialogues, 20 million utterances, and 150 million words, which contains conversations about after-sales topics between users and customer service staffs in an e-commerce scenario. JDDC was updated with multi-modal customer service information in 2021 [zhao2021jddc].
  • E-commerce dialogue corpus dataset [zhang2018modeling]: The dataset is collected from the real-world conversations between customers and customer service staff on Taobao. It contains over five types of conversations (i.e., commodity consultation, logistics express, recommendation, negotiation, and chitchat) based on over 20 commodities. [^11] https://drive.google.com/file/d/154J-neBo20ABtSmJDvm7DK0eTuieAuvw/view?usp=sharing
Table A.5 summarizes the key statistics for the datasets listed above.
@@
JDDC e-commerce dialogue dataset 1,024,196 20,451,337 150,716,172 7.4 20 [chen2020jddc]